OpenStudy (bossimbacon):

which graph is the graph of the system of equations? x + 2y = 2 4x - y = 4 @pooja195

pooja195 (pooja195):

The question is missing info.

OpenStudy (bossimbacon):

there is a bunch of graphs but dont know how to show it.

pooja195 (pooja195):

scrnshot them

OpenStudy (bossimbacon):

but if we complete this equation then we can graph it and i can choose the right ansswer

pooja195 (pooja195):

ok, isolate x x + 2y = 2

OpenStudy (bossimbacon):

subtract it from both sides 2y = x-2

pooja195 (pooja195):

x=2-2y

pooja195 (pooja195):

x=2-2y <--e1 4x -y=4 <--e2 Take equation 1 and plug it into e2 4(2-2y)-y=4 solve for y

OpenStudy (bossimbacon):

8 - 8y - y = 4

OpenStudy (bossimbacon):

-8 - y?

pooja195 (pooja195):

4x−y=4 4(−2y+2)−y=4 −9y+8=4(Simplify both sides of the equation) −9y+8+−8=4+−8(Add -8 to both sides) −9y=−4 divide both sides by -9 y=?

OpenStudy (bossimbacon):

2.25?

pooja195 (pooja195):

keep as fraction 4/9

pooja195 (pooja195):

Now take y=4/9 and plug it into e1 x=−2(4/9)+2 Solve for x

OpenStudy (bossimbacon):

2 * 4/9?

pooja195 (pooja195):

* -2 * 4/9

OpenStudy (bossimbacon):

the answer is 1.11111111

pooja195 (pooja195):

as the final answer?

OpenStudy (bossimbacon):

yeah :/

pooja195 (pooja195):

Correct keep it in fraction form 10/9 so we have x=10/9 y=4/9

OpenStudy (bossimbacon):

so on a graph how would it look like? @pooja195

pooja195 (pooja195):

Idk, you didn't post the graphs.

pooja195 (pooja195):

https://www.desmos.com/calculator type your equations in like this http://prntscr.com/9om0g1 And get your answer. That's all i can tell you to do since the graphs were not posted.

OpenStudy (bossimbacon):

thank you so much again! not many people have the patience to help me throughout the entire thing so thank you!

pooja195 (pooja195):

yw ^_^ We are all here to learn! :D